[Hallc_running] [New Logentry] RC: Daily update

mdiefent at jlab.org mdiefent at jlab.org
Tue Apr 17 12:25:01 EDT 2018

Logentry Text:
<p><strong>Update on day and swing shifts of April 16 and owl shift of April 17 </strong></p><p><strong>CEBAF status</strong>:</p><ul><li>Beam studies during the day shift (https://logbooks.jlab.org/entry/3561759). 0.8 hours of beam time. </li><li>Beam studies during the swing shift (https://logbooks.jlab.org/entry/3561978). Currents up to 80muA (https://logbooks.jlab.org/entry/3561773). 3.0 hours of beam time. </li><li>Stable beam during the owl shift (https://logbooks.jlab.org/entry/3562172). 7.1 hours of beam time. RF recovery starting on April 17 at 8:30 a.m..</li></ul><p><strong>Hall C status</strong>:</p><ul><li>Hall C is configured for the SIDIS experiment (E12-09-017). The HMS is setup for detecting electrons and the SHMS is setup for detecting charged pions and kaons. Details on the experimental setup are available on: https://hallcweb.jlab.org/wiki/index.php/E12-09-017_Run_Plan#Experimental_setup </li><li>For various theta angles in the HMS, the beam current 
 is reduced significantly to keep the pTRIG1 rate below 500kHz. </li><li>Controlled access during the beam studies to work on the problems with: </li><ul><li>the HMS Dipole power supply (https://logbooks.jlab.org/entry/3561651) and </li><li>the HMS Q3 JT8 (https://logbooks.jlab.org/entry/3561662). </li></ul></ul><p><strong>Data taking</strong>:</p><ul><li>The run plan has been updated on April 16 to optimize the physics reach. The detailed plan is available on: https://hallcweb.jlab.org/wiki/index.php/E12-09-017_Run_Plan </li><li>For kinematic setting 1, we have completed 10 out of 14 measurements with positive SHMS polarimetry and also 10 out of 14 measurements with negative SHMS polarimetry.</li><li>Harp scans and runs for the BCM calibration for low beam currents (4muA - 10 muA) during swing shift (https://logbooks.jlab.org/entry/3561869). </li><li>Very successful owl shift by A. Bandari, A. Karki, and C. Yero with 6.4 hours of ABU (https://logbooks.j
 lab.org/entry/3562004).</li><li>The data tak!
 ing of the SIDIS experiment with various configuration changes between the runs requires three persons on shift. We have many open slots in the shift schedule (https://misportal.jlab.org/mis/physics/shiftSchedule/?experimentRunId=HALLC-2017-2018), in particular for the coming weekend and the coming week. We need more volunteers for shifts.</li></ul><p><strong>Hall C Daily RC meeting</strong>: We will meet today at 1:30 p.m. in the conference room located on the second floor of the counting house. Remote participation will be possible via Blue Jeans (Meeting ID: 543903672). You can join via:</p><ul><li>browser: https://bluejeans.com/543903672 </li><li>phone: (888) 240-2560 (US toll free), list of international numbers: https://bluejeans.com/numbers</li></ul>


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3562325
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20180417/2d34aa11/attachment.html>

More information about the Hallc_running mailing list